OpenStudy (anonymous):

Let f(x) = −4x + 7 and g(x) = 10x − 6. Find f(g(x)). −40x + 24 −40x + 31 −40x + 64 −40x + 70

hero (hero):

1. First write the equation for f(x) f(x) = -4x + 7 2. Replace x with g(x) on both sides: f(g(x)) = -4(g(x)) + 7 3. On the right side only, replace g(x) with 10x - 6: f(g(x)) = -4(10x - 6) + 7 4. Next, simplify the right side.

hero (hero):

@Player_2 can you do step 4?

OpenStudy (anonymous):

i would distube right?

hero (hero):

Yes, distribution is necessary to simplify the expression. Go ahead and give it a try, then post your result here.

OpenStudy (anonymous):

f(g(x))=-40x+24+7

hero (hero):

Add 24 + 7 to finish simplifying.

OpenStudy (anonymous):

31

OpenStudy (anonymous):

-40x+31

hero (hero):

Correct. Great job.

OpenStudy (anonymous):

thanks for the help @Hero

hero (hero):

yw